OpenStudy (anonymous):

8 X 10^(2x-7) = 3 Need help solving the exponential/logarithmic function

OpenStudy (anonymous):

divide both sides by 8, then take the base ten log of both sides... what do u get as a result?

OpenStudy (kirbykirby):

\[ 8 \times 10^{2x-7} = 3\\ 10^{2x-7}=\frac{3}{8}\\ \log(10^{2x-7})=\frac{3}{8}\\ (2x-7)\log10=\log\left( \frac{3}{8}\right)\\ 2x - 7 = \log\left( \frac{3}{8}\right), \text{since }\log10=1\\ 2x = \log\left( \frac{3}{8}\right)+7\\ x = \frac{\log\left( \frac{3}{8}\right)+7}{2}\]